OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 6y = –15 x – 3y = –3

OpenStudy (anonymous):

Equation 1: x-6y=-15 Solving for x yields: x=-15+6y (we added +6y to both left and right hand sides) Now if we substitute this into Equation 2: x-3y=-3 you have (-15+6y) -3y=-3 (because x=-15+6y Simplifying, you get -15+3y=-3 can you take it from here?